OpenStudy (board_man):

If y varies directly as x, and y = 15 as x = -3, find y for the x-value of 4.

OpenStudy (welshfella):

y = kx where k is the constant of variation the first step is to find the value of k by substituting values for x and y. These are x = -3 and y = 15. Can you do that?

OpenStudy (welshfella):

y = kx is direct variation.

OpenStudy (board_man):

would it be.... x=-20?

OpenStudy (welshfella):

I'll start you off y = k x 15 = k* -3 tofind k we divide both sides of the equation by -3.

OpenStudy (board_man):

-3/15=4/x is the first step right?

OpenStudy (queenie):

if it varies directly, u can set it up as a proportion

OpenStudy (welshfella):

yes - you can do it that way

OpenStudy (welshfella):

-20 is correct

OpenStudy (board_man):

I wish I could give you more medals :(

OpenStudy (welshfella):

thats OK yw

OpenStudy (welshfella):

not that k is -5 so if x = 4 y = -5*4 = -20

OpenStudy (board_man):

You're awesome! no wonder you're the honorary professor